What was found
F/A'S AND PAX COMPLAINED ABOUT CABIN TEMP. MOSTLY STAYING TOO COLD WITH TEMP SELECTORS IN 2 O`CLOCK POSITION AND THOUGH FURTHER MOVEMENT TOWARD 3 O`CLOCK RESULTED IN TEMP BEING VERY HOT. ACFT GROUNDED. FOUND FWD AND AFT CABIN TEMP SENSOR FILTERS CLOGGED. CLEANED FILTERS REF AMM 21-61-06-100/801. OPS CKS OF BOTH PACKS GOOD. PERFORMED BITE CK OF BOTH CONTROLLERS REF FIM 21-62, TASK 801, NO FAULTS. GROUND FUNCTIONAL TEST PASSED.